I have no idea do even start this problem I reallly need help.
The album thriller was released in december of 1983, when the king of pop was worth $37 million. when thriller finally fell from the billboard top 10 in july 1985 micheal's net worth was $193 million.
Using december 1983 as a reference month, determine a liner function that can ve used to estimate Micheals net worth during the thriller albums billboard top 10 period. let your time variable represent the number of months since december.
I was thinking that i would have to do something like subtract 37 from 193 or something but im not confident about it..help?!?!

Answer by stanbon(75887)

The album thriller was released in december of 1983, when the king of pop was worth $37 million. when thriller finally fell from the billboard top 10 in july 1985 micheal's net worth was $193 million.
Using december 1983 as a reference month, determine a liner function that can ve used to estimate Micheals net worth during the thriller albums billboard top 10 period. let your time variable represent the number of months since december.
--------------------
You have two points (0,37) ; 0 represents december, 1983
and (19,193) ; 19 represents july, 1985
---------------------------------------------
slope = (193-37)/19 = (156)/19
intercept occurs when x = 0 ; intercept is 37
---------------------------------------------
Equation:
worth(in millions) = (156/19)(time variable) + 37
W = (156/19)t + 37
========================================================
Cheers,
Stan H.